Find the lengths of the sides of a triangle whose perimeter is 37. The measure of the first side of the triangle is 8 less than the second side, and the secoond side is twice the length of the third.

Answer by cleomenius(959)

A = 2x - 8
B = 2x
C = x
====================
2x - 8 + 2x + x = 37
5x = 45
x = 9
=====================
A = 18 - 8 = 10
B = 18
C = 9
10 + 18 + 9 = 37
This does check.
cleomenius